HUBBLE SPACE TELESCOPE DAILY REPORT #2252 PERIOD COVERED: 0000Z (UTC) 11/16/98 - 0000Z (UTC) 11/17/98 Daily Status Report as of 321/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Four Sets of WF/PC-2 6731 (Exploring PN Production as a Probe of Elliptical Galaxy Stellar Populations) The WF/PC-2 was used to make observations of the galaxy NGC-4278 in order to explore planetary nebulae production as a function of elliptical galaxy stellar population. In addition to being an excellent distance indicator, the O III Lambda 5007 A0 planetary nebula luminosity function (PNLF) is potentially a powerful probe of stellar population within these galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Nine Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.4 Completed Two Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.5 Completed STIS/MA2 7650 (STIS Yearly CCD Wavelength Monitor) The Space Telescope Imaging Spectrograph (MA2) was used to perform its yearly wavelength calibration monitoring program. Wavelength dispersion solutions will be determined on a yearly basis as part of a long-term monitoring program. The observations were completed as planned, and no anomalies were reported. 1.6 Completed Five Sets of NIC/1/2/3 7961 (Flats: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. These flats will provide a monitor for particulate comtamination and detector lateral position {from the coronagraphic spot and FDA vignetting}. The proposal completed with no reported problems. 1.7 Completed Three Sets of STIS/MA1 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.8 Completed Five Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed Two Sets of FOC/96 6689 (Determination of the Distances and Masses of Two Galactic Cepheids) The Faint Object Camera (f/96) was used to make observations of galactic Cepheid star HD183344. The angular separation of its binary components will be measured from the images. When combined with the information from the spectrographic orbits, the data will yield the masses and distances to the Cepheid. AN internal flat calibration observation was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Three Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.11 Completed NIC/2/3 7962 (NICMOS Warm-up Focus Monitoring) The Near Infrared Camera (Cameras Two and Three) was used to make observations of the globular cluster and nebula of NGC-3603 as part of the NICMOS warm-up focus monitoring. This proposal is used to determine and monitor the optimal focus settings for NIC/2 and NIC/3 during the dewar warm-up period at the end of 1998. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 6698 (High Resolution Imaging of Bubble and Superbubbles in H II Regions) The WF/PC-2 was used to make high resolution observations of the H II region N51D. Massive stars dynamically interact with their ambient interstellar medium via fast stellar winds during their lifetime and via supernova blasts at their demise. The dynamical interaction causes the physical structure of an H II region to evolve along with its ionizing massive stars. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 7336 (Cosmological Parameters Omega and Lambda From High-Redshift Type Ia Supernovae) The WF/PC-2 was used to make observations of the high redshifted Type Ia supernova 5H1.3ZA. HST can make a direct measurement of the cosmological parameters, Omega, Lambda, and curvature, using Type Ia supernovae as calibrated "standard candles." The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed WF/PC-2 6347 (A Search for Jets in Planetary Nebulae) The WF/PC-2 was used to make observations of the planetary nebulae (class PNe) NGC-2392 in order to search for the presence of possible narrow jets. The current model for PNe formation indicates that jets are an unexpected phenomena. Yet there is a growing body of circumstantial evidence that suggests narrow jets are common feature in PNe. The observations were completed as scheduled, and no anomalies were noted. 1.15 Completed STIS/CCD/MA2 7317 (Spatial Distribution of Ozone and Sulphur Dioxide on Ganymede and Europa) The Space Telescope Imaging Telescope (CCD/ MA2) was used to investigate the O_2 and O_3 in spectra of Ganymede's trailing hemisphere. Ganymede's icy surface is an unexpectedly complex and dynamic environment. Europa, exposed to a slightly different radiation environment, has SO_2 on its trailing hemisphere with no trace of either O_3 or O_2. The new capability to measurespectra at 0.1 arcsec spatial scales offered by STIS gives us the opportunity to address such questions by mapping the distribution of O_3 and OH on Ganymede and SO_2 on Europa. The proposal completed as planned, with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 10 Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 25 Successful : 25 2.3 Operations Notes: The Leonid Meteor Shower (i.e., the Temple-Tuttle Wake) is expected to peak during a window of 321/1640Z through 2243Z. During the identified "storm protection" window at 321/1443Z through 322/0043Z, the HST attitude will be maintained in inertial pointing at RA 330.65 degrees and DEC -23.55 degrees (-V3 towards the storm's radiant) to minimize the vehicle's cross sectional area which is facing the storm radiant. The Solar arrays will also be maintained in a fixed ordination, parallel with the V1 axis. All command activities will be minimized during the "protection" window. Stored command uplinks and recorder playbacks have been replanned to avoid this window. There were three IP Data hits at 320/1150Z, 1226Z and 1238Z. TTR 20382 was written. The SSR EDAC Error counter was cleared four times (ROP SR-1) 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH